OpenStudy (anonymous):

3 moles of an ideal monoatomic gas at 27 C and 1 atm pressure are subjected to a reversible adiabatic Compression to half the initial volume calculate Q , W and U in calories?

OpenStudy (dls):

U=5400J?

OpenStudy (anonymous):

Nope...)

OpenStudy (anonymous):

\[\Delta U = Q + W\] Here Q=0 Since Adiabatic Process So \[\Delta U = W\]

OpenStudy (anonymous):

\[\frac{ T1 }{ T2}=(\frac{ V2 }{ v1 })^{\gamma -1}\]

OpenStudy (kainui):

This problem looks like you'll need to use gamma, do you remember the relation?

OpenStudy (kainui):

Yes exactly.

OpenStudy (kainui):

Now go through the fun of converting everything around to SI units and then calories and other pointless nonsense lol.

OpenStudy (anonymous):

dW = PdV You can find V from PV=nRT, I believe.

OpenStudy (kainui):

You can use that to calculate the initial volume.

OpenStudy (anonymous):

May be Finding T2 and Solving Will be Easy...i think..

OpenStudy (anonymous):

V_2 = 0.5V_1

OpenStudy (kainui):

U=CvdT right? I think you have a way of finding Cv for an ideal gas, just look above and then you have a change in temperature to plug in after the "integration" so you really don't need to go very far.

OpenStudy (anonymous):